# Enhancing Image Layout Control with Loss-Guided Diffusion Models

### v1

@@ A.2 Score Matching
Since $q(x_{t}|x_{0})$ is a normal distribution, we know that
@@
for some $\lambda(t)$, where
@@
It is known that this loss is minimized when $\mathbf{s_{\theta}}(\mathbf{x}_{t},t)=\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t})$ Vincent (2011), so given enough parameters, $\mathbf{s_{\theta}}(\mathbf{x}_{t},t)$ will converge to $\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t})$ almost everywhere. With this, the mean $\mu_{\theta}(\mathbf{x}_{t},t)$ of the reverse process can be understood as
$\mu_{\theta}(\mathbf{x}_{t},t)=\sqrt{\bar{\alpha}_{t-1}}D_{\theta}(\mathbf{x}_{t},t)\approx\mathbb{E}_{q}[\sqrt{\bar{\alpha}_{t-1}}\mathbf{x}_{0}|\mathbf{x}_{t}],$ (24)
@@
$D_{\theta}(\mathbf{x}_{t},t):=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x}_{t}+\beta_{t}s_{\theta}(\mathbf{x}_{t},t))$ (25)
@@
$\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x}_{t}+\beta_{t}\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}))=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\mathbb{E}_{q}[\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}|\mathbf{x}_{t}]=\mathbb{E}_{q}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ (26)
Efron (2011). Given an approximation to the score function, it is possible to sample from $p_{\theta}(\mathbf{x}_{0})$ using annealed Langevin dynamics Song and Ermon (2019).
@@ A.3 Stochastic Differential Equations
Song et al. (2021) showed that the forward process of DDPM can viewed as a discretization of the stochastic differential equation (SDE)
@@
Appendix B Additional Experiments
We provide two additional sets of comparisons between our proposed method (iLGD), BoxDiff, and Stable Diffusion. In Figure B.1, we compare the three methods using same prompts and bounding boxes as in Figure 4, but using a different random seed for each set of images. In Figure B.1, we compare the methods using an entirely new set of prompts and bounding boxes.
Figure B.1: A comparison of iLGD against BoxDiff and Stable Diffusion, using the same prompts as Figure 4 but different random seeds, with the seed kept the same across each set of images.
Figure B.2: A comparison of iLGD against BoxDiff and Stable Diffusion, for an entirely new set of prompts. The random seed kept the same across each set of images.
Appendix C Detailed Methods
@@
In this section, we describe in detail how we obtain the AP@50 scores in Table 2. In classical object detection, a model is trained to detect and localize objects of certain classes in an image, typically by predicting a bounding box which fully encloses the object. The accuracy of the model’s predicted bounding box, $B_{\text{p}}$, is evaluated by comparison to the corresponding ground truth bounding box, $B_{\text{gt}}$. More specifically, we compute the intersection over union (IOU) over the pair of bounding boxes:
$\text{IOU}=\frac{\text{area}(B_{\text{p}}\cap B_{\text{gt}})}{\text{area}(B_{\text{p}}\cup B_{\text{gt}})}.$ (28)
The IOU is then compared to a threshold $t$, such that, if $\text{IOU}\geq t$, then the detection is classified as correct. If not, then the detection is classified as incorrect. In our case, we follow Li et al. (2021) and treat the object detection model as an oracle, where we assume that it provides the bounding boxes of objects in a given image with perfect accuracy, and use the desired layout in place of the ground truth bounding boxes. We generate an image according to our desired layout and apply the object detection model to the generated image to obtain a set of predicted bounding boxes. Finally, to evaluate how similar the layout of the generated image is to the desired layout, we compare each predicted bounding box, $B_{\text{p}}$, to the corresponding ground truth bounding box, $B_{\text{gt}}$, by computing their IOU. We use a IOU threshold of 0.5.
@@
Wang et al. (2023) suggest using the pair {good photo, bad photo} instead of {high quality, low quality} to measure quality, as they find that it corresponds better to human preferences. However, we choose the latter to remain agnostic to the image’s style, as we believe the former carries with it a stylistic bias, due to the word “photo.”

### v2

@@ A.2 Score Matching
Since $q(\mathbf{x}_{t}|\mathbf{x}_{0})$ is a normal distribution, we know that
@@
for some positive function $\lambda(t)$, where
@@
It is known that this loss is minimized when $\mathbf{s_{\theta}}(\mathbf{x}_{t},t)=\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t})$ Vincent (2011), so given enough parameters, $\mathbf{s_{\theta}}(\mathbf{x}_{t},t)$ will converge to $\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t})$ almost everywhere. Given an approximation to the score function, it is possible to sample from $p_{\theta}(\mathbf{x}_{0})$ using annealed Langevin dynamics Song and Ermon (2019).
Letting the forward process posterior mean $\tilde{\mu}_{t}$ be defined by $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})\sim\mathcal{N}(\mathbf{x}_{t-1};\tilde{\mu}_{t}(\mathbf{x}_{t},\mathbf{x}_{0}),\tilde{\beta}_{t}\mathbf{I})$, we have that
$\tilde{\mu}_{t}(\mathbf{x}_{t},\mathbf{x}_{0}):=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\mathbf{x}_{0}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\mathbf{x}_{t}$ (18)
(see Ho et al. (2020)). With this, the mean $\mu_{\theta}(\mathbf{x}_{t},t)$ of the reverse process can be understood as
$\mu_{\theta}(\mathbf{x}_{t},t)=\tilde{\mu}_{t}(\mathbf{x}_{t},D_{\theta}(\mathbf{x}_{t},t)),$ (19)
@@
$D_{\theta}(\mathbf{x}_{t},t):=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x}_{t}+(1-\bar{\alpha}_{t})s_{\theta}(\mathbf{x}_{t},t))$ (20)
@@
$\begin{split}&\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{x}_{t}+(1-\bar{\alpha}_{t})\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}))\\ &\hskip 30.00005pt=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\mathbb{E}_{q}[\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}|\mathbf{x}_{t}]=\mathbb{E}_{q}[\mathbf{x}_{0}|\mathbf{x}_{t}]\end{split}$ (21)
(see Efron (2011)).
@@ A.3 Stochastic Differential Equations
Song et al. Song et al. (2021) showed that the forward process of DDPM can viewed as a discretization of the stochastic differential equation (SDE)
@@
A.4 Classifier-Free Guidance
In order to generate images following a user-supplied text prompt, the denoiser $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{y})$ of a latent diffusion model is trained with an additional input given by a sequence of token embeddings $\mathbf{y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{k}\}$. A single denoiser, usually a UNet, is trained over a variety of text prompts, and the token embeddings influence the denoiser by a cross-attention mechanism in both the contractive and expansive layers. Ho and Salimans Ho and Salimans (2021) found that, rather than sampling images using the conditional denoiser alone, better results can be obtained by taking a combination of conditional and unconditional noise estimates,
Appendix B Detailed Methods
Implementation Details
We implement our method on the official Stable Diffusion v1.4 model Rombach et al. (2022) from HuggingFace. All images are generated using 50 denoising steps and a classifier-free guidance scale of 7.5, unless otherwise noted. We use the noise scheduler LMSDiscreteScheduler Karras et al. (2022) provided by HuggingFace. Experiments are conducted on an NVIDIA TESLA V100 GPU.
When attempting to control the layout of a generated image, we find that skipping the first step, so that it remains a standard denoising step, leads to better results. We do this for all experiments conducted in this paper which use either injection or loss guidance or both. In iLGD, we use $\eta=0.48$, $\nu^{\prime}=0.75$, $t_{\text{loss}}=25$, and $t_{\text{inject}}=10$, unless otherwise noted. In our ablation experiments, we keep the injection strength at $\nu^{\prime}=0.75$ when performing just attention injection. When performing just loss guidance, we increase the loss-guidance strength to $\eta=1$, in order to make loss guidance alone exert sufficient influence over the final image layouts.
In our comparisons with BoxDiff, we maintain the default parameters the authors provide in their implementation. We start with $\alpha_{T}=20$, which decays linearly to $\alpha_{0}=10$, and perform guidance for 25 iterations out of a total of 50 denoising steps.
In our comparisons with the method of Chen et al., we also maintain the default parameters the authors provide in their implementation, setting the loss scale factor to $\eta=30$.
@@ Appendix B Detailed Methods
In this section, we describe in detail how we obtain the AP@50 scores in Table 2. In classical object detection, a model is trained to detect and localize objects of certain classes in an image, typically by predicting a bounding box which fully encloses the object. The accuracy of the model’s predicted bounding box, $B_{p}$, is evaluated by comparison to the corresponding ground truth bounding box, $B_{gt}$. More specifically, we compute the intersection over union (IOU) over the pair of bounding boxes:
$\text{IOU}=\frac{\text{area}(B_{p}\cap B_{gt})}{\text{area}(B_{p}\cup B_{gt})}.$ (24)
The IOU is then compared to a threshold $t$, such that, if $\text{IOU}\geq t$, then the detection is classified as correct. If not, then the detection is classified as incorrect. In our case, we follow Li et al. Li et al. (2021) and treat the object detection model as an oracle, where we assume that it provides the bounding boxes of objects in a given image with perfect accuracy. In particular, we first define a layout through a set of ground truth bounding boxes, describing the desired positions of each object. We then generate an image according to this layout, and subsequently apply the object detection model to the generated image to obtain a set of predicted bounding boxes. Finally, to evaluate how similar the layout of the generated image is to the desired layout, we compare each predicted bounding box, $B_{p}$, to the corresponding ground truth bounding box, $B_{gt}$, by computing their IOU. We use a IOU threshold of 0.5.
@@
Wang et al. Wang et al. (2023) suggest using the pair {good photo, bad photo} instead of {high quality, low quality} to measure quality, as they find that it corresponds better to human preferences. However, we choose the latter to remain agnostic to the image’s style, as we believe the former carries with it a stylistic bias, due to the word “photo.”
@@
Appendix C Additional Experiments
We provide two additional sets of comparisons between our proposed method (iLGD), BoxDiff Xie et al. (2023), Chen et al. Chen et al. (2024), MultiDiffusion Bar-Tal et al. (2023), and Stable Diffusion Rombach et al. (2022) . In Figure 1, we compare the three methods using same prompts and bounding boxes as in Figure 3, but using a different random seed for each set of images. In Figure 2, we compare the methods using an entirely new set of prompts and bounding boxes.
Figure 1: A comparison of iLGD against BoxDiff, Chen et al., MultiDiffusion, and Stable Diffusion, using the same prompts as Figure 3 but different random seeds, with the seed kept the same across each set of images.
Figure 2: A comparison of iLGD against BoxDiff, Chen et al., MultiDiffusion, and Stable Diffusion. The random seed kept the same across each set of images.